Determine the force of gravity of a person if the distance is 10 times from Earth's center and when the force of gravity of an average person on Earth's surface is about 700 N.

1
Expert's answer
2022-01-03T10:46:03-0500
"F_{new}=\dfrac{Gm_pM_E}{(10d)^2}=\dfrac{Gm_pM_E}{100d^2}=\dfrac{F}{100}=\dfrac{700\ N}{100}=7\ N."